# Tiling Spaces and the Expanding Universe: Bridging Quantum Mechanics and Cosmology

### v1

@@ Abstract
We propose a heuristic model of the universe as a growing quasicrystal projected from a higher-dimensional lattice. By extending the Schrödinger equation for a particle in a box with time-dependent boundaries, we derive an equation that resembles the Friedmann equation, addressing the Hubble tension. This model incorporates phonons and phasons, providing insights into cosmic-scale dynamics and the universe’s expansion. We outline a pre-inflation tiling space phase with quantum error correction, a radiation phase dominated by quasiparticles, and a post-radiation phase with the emergence of matter. Our hypothesis, which posits that the universe is a growing quasicrystal, suggests that the necessity for an inflationary period may be obviated. Furthermore, phonon arising from this quasicrystalline structure could act as dark matter, influencing the dynamics of ordinary matter while remaining mostly undetectable by electromagnetic interactions. This hypothesis draws parallels with other crystalline matter at atomic scales that are fundamentally quantum in nature. We also explore how the notion of tiling space can support continuous symmetry atop a discrete structure, providing a novel framework for understanding the universe’s expansion and underlying structure. Consequently, it is logical to start with quantum mechanics as the foundation of our model. Further development could enhance our understanding of cosmic expansion and the underlying structure of the universe.
Keywords Quantum Gravity  $\cdot$ Cosmology  $\cdot$ Hubble Tension  $\cdot$ Quasicrystals
@@ 1 Introduction
Recent developments in theoretical physics have propelled us toward a deeper understanding of spacetime and gravity. The field of emergent gravity [1] suggests that spacetime and gravity might be the byproducts of fundamental quantum mechanical processes. This paradigm shift is supported by theories such as loop quantum gravity and spin foam [2], as well as superstring theory and M theory [3], all of which propose a quantum-level discreteness underlying gravity.
A key focus since the advent of perturbative superstring theory has been the quest to unravel quantum gravity [4]. This endeavor involves comprehending the theory’s foundational principles and extracting insights about Planck-scale physics [5]. Consequently, there is a growing trend to incorporate concepts from condensed matter or solid-state physics, such as viewing spacetime as a deformable solid or crystal [6]. More recently, quasicrystals have been investigated [7, 8].
@@
Our research adopts a dual approach within this context. First, we start with conventional quantum mechanics. While the universe’s expansion is typically addressed from a classical perspective on large scales, this expansion must also extend down to microscopic scales, as suggested by various approaches to quantum cosmology [12, 13]. Unlike conventional quantum cosmology, we analyze the wave function of a particle with mass $m$ in a specific spatial domain using the Schrödinger equation, thereby focusing on intermediate scales. Crucially, we posit that the wavenumber, wavelength, and quantum amplitudes become functions of the dynamical size of the box and are subject to the Hubble law explicitly. Considerations about the dynamic boundary of the spatial domain lead us to a version of the Friedmann equation, crucial in cosmological models, thereby underscoring the emergence of classical spacetime from quantum mechanics.
@@
Second, drawing inspiration from the concept of a deformable solid or crystal and quantum cosmology, which deals with discreteness, we propose modeling spacetime as a growing quasicrystal. Characterized by long-range order without periodicity, this structure reflects the aperiodic patterns found in condensed matter physics. Our principal finding suggests that the cosmological scale factor might be interpreted as an inherent scaling characteristic within a quasicrystal tiling space framework. We also consider the implications of this model for the Hubble tension, offering a fresh perspective on the general evolution of the universe and providing new insights into other issues in current cosmology models, such as dark matter and dark energy.
@@
Our research adopts a dual approach that intersects at the idea of rescaling the underlying geometry, accompanied by emerging constraints pertinent to physics at varying scales. This approach aligns with our methodology of conceptualizing spacetime through the lens of a quasicrystalline structure. By integrating scale invariance into our framework, we aim to forge a connection between the microscale realms of quantum mechanics and the macroscale phenomena observed in cosmology. This synthesis offers a comprehensive and unified perspective, potentially enhancing our understanding of the universe’s nature.
@@
In the following sections, we will further elaborate on the mathematical foundations of our model, exploring how energy considerations translate into cosmological parameters. Moreover, we will delve into the concept of spacetime as a tiling space, drawing parallels with quasiperiodic structures in condensed matter physics to bolster our hypothesis of emergent spacetime and gravity.
2 From Schrödinger to Friedmann Equation
$G_{\mu\nu}+\Lambda g_{\mu\nu}=\frac{8\pi G}{c^{4}}T_{\mu\nu},$ (1)
where $G_{\mu\nu}$ is the Einstein tensor representing the curvature of spacetime, $g_{\mu\nu}$ is the metric tensor describing the geometry of spacetime, $T_{\mu\nu}$ is the stress–energy tensor reflecting the distribution and flow of energy and momentum in spacetime, $\Lambda$ is the cosmological constant, $G$ is the Newtonian constant of gravitation, and $c$ is the speed of light in vacuum. On the other hand, quantum mechanics introduces the principle of wave-particle duality, which manifests both wave-like and particle-like properties in matter and energy, as illustrated by the De Broglie relations and the Schrödinger equation-Equations 2, 3. In this sense, both equations of general relativity and quantum mechanics have matter on one side and something else on the other side of an equality. This “something else” being geometry-like or wave-like. In this discourse, we explore a simplified model within the context of cosmology, where the wave aspect of quantum mechanics is proportionally linked to the geometry of spacetime. This approach specifically considers the cosmological scale factor that characterizes the rescaling of spacetime within a homogeneous and isotropic universe described by a perfect fluid.
The Friedmann equations, which are central to our understanding of the universe’s expansion, are primarily derived from general relativity and frameworks that integrates classical mechanics and thermodynamics. This exploration contemplates whether analogous dynamics might emerge from applying space rescaling concepts to a local particle quantum state. The conventional understanding posits that the universe expands on large scales, particularly in the vast regions between galaxies, while gravitational forces within galaxies maintain their structural cohesion. However, this view does not necessarily preclude the possibility of spacetime expansion within matter-dominated regions, which may be challenging to observe directly.
Given that matter fundamentally are quantum phenomena and acknowledging the absence of a unified quantum gravity theory, it is plausible to infer that the cosmological expansion observed at macroscopic scales originates from the spacetime fabric at the microscopic level, possibly extending down to the Planck scale. While general relativity provides a framework for understanding how matter influences spacetime curvature, the ability of classical gravity to keep matter bound at macroscopic scales, even amidst the potential expansion of the underlying spacetime, raises intriguing questions about the emergence of matter from fundamental quantum dynamics.
To explore these ideas further, our focus shifts to examining the behavior of an existing quantum state within the context of universal expansion. By investigating how quantum states respond to changes in the scale of spacetime, we may gain new insights into the interplay between quantum mechanics and cosmology, potentially shedding light on the origin of the observed cosmological expansion and its relationship to the microscopic structure of spacetime.
Consider the time-independent Schrödinger Equation for the wave function (energy eigenfunction) $\Psi(x)$, with $x$ representing $(x,y,z)$, of a particle with mass $m$ under potential $V(x)$. The equation is given by:
$\nabla^{2}\Psi=-\frac{2m(E-V)}{\hbar^{2}}\Psi\quad,$ (2)
For a particle under a potential $V(x)$, the time-independent Schrödinger equation can yield different forms of solutions depending on the relationship between the energy $E$ of the particle and the potential $V$. We consider the particle in a box situation with a finite potential inside and infinite outside. The box may be taken to be spherical with a given radius. Specifically, inside the box, for a constant potential, being zero, or when $E>V$, the solutions are of the form of plane waves. Conversely, for $E<V$, the solutions differ significantly. When these solutions are substituted back into the time-independent Schrödinger equation, we obtain:
$k^{2}=\frac{2m(E-V)}{\hbar^{2}},$ (3)
where $k$ is the wavenumber.
$\eta(t_{2})=\frac{a(t_{2})^{3}}{a(t_{1})^{3}}\eta(t_{1}),$ (4)
Introducing time is essential to account for the rescaling effects. The problem now should be addressed by the time-dependent Schrödinger equation, as the potential depends on the dynamic radius or volume. This problem has been studied using various approaches [14] and references therein. In this context, the energies become dynamic, and a notion of energy density in quantum mechanics arises naturally [15]. Crucially, the wavenumber, wavelength, and quantum amplitudes become functions of the dynamical size of the box, and we posit that they are subject to the Hubble law explicitly. By connecting the geometry with the wave aspect of the particle-wave duality mentioned earlier, we aim to derive Friedmann-like equations.
Revisiting equation (3), we modify it by dividing both sides by the dynamical volume $\eta$, which confines the particle in space. Equation (3) encapsulates the wave-particle duality, highlighting that the matter wavelength, $\lambda=\frac{2\pi}{k}$, is a fundamental quantum mechanical attribute that determines the probability density of locating the particle within the configuration space. Notably, a particle’s wavelength, inversely related to its momentum (and thus energy), serves as a bridge to cosmological dynamics. By invoking the Hubble law, expressed as $v=H\lambda$, where $H=\frac{\dot{a}}{a}$ is the time-dependent Hubble parameter, we correlate the particle’s wavelength to the cosmic scale factor $a(t)$ while keeping the energy constant. This association suggests an emergent classical kinetic energy tied to the velocity $v$ associated to the expansion. Consequently, equation (3) is transformed into
where $K=\frac{1}{2}mv^{2}$ is the kinetic energy, $\rho=E/\eta$ represents the energy density, and setting $V=0$. Note that we are simplifying the concept of energy density here. The particle energy is time-dependent, and any particle would be subject to the same expansion that occurs everywhere. This brings a notion of a particle horizon and an energy density within each particle volume expansion. This equation bears a striking resemblance to the first Friedmann equation, suggesting that the dynamics of a quantum particle in a box with a time-dependent boundary may provide insights into the expansion of the universe. By establishing a connection between the particle’s wave nature and the cosmic scale factor. To align with the usual first Friedmann equation’s structure, the term $\frac{K\eta}{\pi^{2}\hbar^{2}}$ must remain constant
$K\eta=\frac{8\pi^{3}\hbar^{2}G}{3c^{2}}$ (6)
$H^{2}=\left(\frac{\dot{a}}{a}\right)^{2}=\frac{8\pi G}{3c^{2}}\rho,$ (7)
$(\hbar kc)^{2}=E^{2}-(mc^{2})^{2}.$ (8)
$C=E-K-V=0,$ (9)
where $E$ is the total energy, $K$ is the kinetic energy, and $V$ is the potential energy, all associated with the dynamical expansion process. This constraint allows us to examine the system’s properties across different scales, connecting the microscopic energy $E$ to macroscopic scales potentially represented by $K$ and $V$. Here, $K$ and $V$ can be viewed as effective parameters encapsulating average or microscopic system properties as the system expands.
The energy-balancing constraint in Equation 9 provides a framework for understanding how the total energy $E$ is partitioned into kinetic and potential components during the expansion process. By considering the conservation of energy and the interplay between these different forms of energy, we can gain insights into the dynamics of the system across various scales. In the context of the relativistic particle in a box with moving walls, the application of this constraint may lead to the emergence of terms analogous to the energy density $\rho$, pressure $p$, and possibly the cosmological constant $\Lambda$ when deriving the relativistic Friedmann-like equation. By incorporating the energy-balancing constraint into the analysis, we can explore how the quantum mechanical properties of the system, such as the wave nature of the particle and the dynamical evolution of the wavefunction, relate to the classical cosmological parameters that describe the expansion of the universe.
Isolating $K$ while keeping $E$ and $V$ independent, we obtain
$E^{2}=K(E+V+V^{2}/K).$ (10)
Since the wave number is inversely proportional to the wavelength, we can apply the Hubble law as done before. Then, Equation 8 can be rewritten as:
with $\beta=v/c$, $\rho=E/\eta$ is the energy density, and $p=V/\eta$ is the pressure. Equation 11 is analog of the Friedmann equation for relativistic matter, derived from the quantum mechanical description of a particle in a box with moving walls. The first term on the right-hand side resembles the classical Friedmann equation. The second term, proportional to $V^{2}$, may be interpreted as related to the cosmological constant $\Lambda$ as it does not explicitly depend on $a$.
To obtain the usual contribution to radiation, we set $mc^{2}=0$. Typically, radiation energy density enters the Friedmann equation as proportional to $1/a^{4}$, while non-relativistic matter is proportional to $1/a^{3}$. Here, we derive a notion of energy density for individual massive particles and photons subject to dynamical boundary conditions related to the expansion of space, and they share a similar form. The $1/a^{4}$ factor for radiation arises from considering the number density of photons, which scales like matter, times the photon energy, which scales with $1/a$ due to redshift as the photon wavelength expands. The redshift here is accounted for through the Hubble law applied to the wavelength.
Importantly, we note that different dispersion relations will lead to different dynamics for the scale parameter, and various relations may be considered. This is in line with the usual approach where different cosmological epochs have different evolutions due to varying energy densities. To motivate our quasicrystalline model presented in the next section, let us consider quantum excitations in lattice and quasicrystalline condensed matter systems. These systems are known for their phonon quasiparticles, which emerge due to the breaking of translational symmetry. Quasicrystals, however, also host phasons—a less understood type of quasiparticle. Phasons are particularly interesting because they are linked to a reduction in dimensions; this occurs as quasicrystals are formed, at least theoretically, from the projection of higher-dimensional lattices onto lower-dimensional spaces.
The dispersion relation for phonons in lattices is similar to the relativistic one in the regime of large wavelengths, but with the speed of light replaced by the speed of sound. In this regime, the phonon dispersion relation is nearly linear, and the crystal can be treated as a continuous elastic medium, with phonons behaving like classical sound waves. However, in the regime of short wavelengths, the dispersion relation becomes nonlinear, and the microscopic details of the lattice become important. In this case, phonons are sensitive to the discrete nature of the lattice, and the nonlinear dispersion relation can lead to phonon dispersion, where phonon packets with different wavelengths propagate at different speeds, causing distortion of the waveform. The transition between these two regimes occurs when the phonon wavelength becomes comparable to the interatomic spacing, with the specific wavevector or wavelength at which this transition happens depending on the details of the crystal structure and the interatomic forces.
In quasicrystals, the situation is more complicated because they do not have a simple primitive cell like periodic crystals but rather a complex, self-similar structure. As a result, the phonon dispersion relations in quasicrystals can exhibit intricate, fractal-like patterns. Despite these differences, the linear dispersion relation for phonons in quasicrystals shares similarities with the phonon branches in periodic crystals, particularly at low frequencies and long wavelengths. We will discuss more about phonons in the next section when discussing the quasicrystalline model, where they arise as candidates to contribute in both radiation (massless phonons) and dark matter regimes (massive phonons).
Phasons in quasicrystals also share similarities with those found in relativistic physics. However, a key difference is the presence of a dispersive imaginary term in the phason dispersion relation, which becomes prominent at very large wavelengths [17, 18]. This imaginary term is related to the diffusive behavior of phasons and is absent in the phonon dispersion relation.
The presence of both phonons and phasons in quasicrystals, along with their distinct dispersion relations, may have important implications for the dynamics of the scale parameter in the quasicrystalline model of the universe. The interplay between these two types of quasiparticles, their different regimes of behavior, and their potential connection to the emergence of cosmological equations could provide new insights into the relationship between quantum mechanics and cosmology.
Using $\omega=E/\hbar$, for phasons, we have
$v^{2}k^{2}=\frac{E^{2}}{\hbar^{2}}+i\gamma\frac{E}{\hbar}$ (12)
where $\gamma$ is the phason friction coefficient. The phason is a goldstone mode that arise under a symmetry breaking mechanism [17]. In the case of lattices, the energy states are defined in a Voronoi cell. For quasicrystals, there are more fundamental cells or tiles, but with one fixed, we can consider the expansion of that cell and, as before, apply the Hubble law to the wavelength. In addition, we can use Equation 10 and multiply by the volume on both sides of Equation 12, yielding
$H^{2}=\frac{K\eta}{4\pi^{2}\hbar^{2}}(\rho+p)+\frac{1}{4\pi^{2}\hbar^{2}}V^{2}+i\gamma\frac{E}{4\pi^{2}\hbar}$ (13)
where we considered the velocity associated to phasons equal to the one of the expansion. To deal with the imaginary part, this equation can be interpreted from the energy relation where in the regime of short wavelength the phason propagation dominates with $v^{2}k^{2}=\frac{E^{2}}{\hbar^{2}}$ leading to a Friedmann like equation similar to Equation 11. In the regime of large wavelength we have $v^{2}k^{2}=i\gamma\frac{E}{\hbar}$ which leads to diffusion and decoupling from expansion contribution. We will come back to these two regimes in context of the quasicrystalline model of the universe in next section.
$\dot{\rho}+3\frac{\dot{a}}{a}(\rho+p)=0.$ (14)
Taking the derivative in Equations (5), (11), and (13), with $V$ constant and $\beta=1$, we get for Equations (5) and (11):
$2H\dot{H}=\alpha\dot{\rho},$ (15)
with $\alpha=K\eta/\pi^{2}\hbar^{2}$. For Equation (13), we have
$2H\dot{H}=\alpha\dot{\rho}+i\gamma\frac{\dot{E}}{4\pi^{2}\hbar}.$ (16)
Now, we can derive the acceleration term in both situations. For Equation (5), substituting the continuity equation into the derivative of the Friedmann-like equation yields (with the original potential equal to zero)
$\frac{\ddot{a}}{a}=-\frac{\alpha}{2}(\rho+3p).$ (17)
For Equations (11), with $m=0$
$\frac{\ddot{a}}{a}=-\frac{\alpha}{2}(\rho+p)+\Lambda$ (18)
with $\Lambda=\frac{1}{\pi^{2}\hbar^{2}}V^{2}$.
For Equation (13), which includes the phason contribution, the acceleration equation becomes
$\frac{\ddot{a}}{a}=-\frac{\alpha}{4}(\rho+p)+\frac{1}{4}\Lambda+\frac{i\gamma}{4\pi^{2}\hbar}(E+\frac{\dot{E}}{2H})$ (19)
The additional term involving the phason friction coefficient $\gamma$ and the time derivative of the energy $\dot{E}$ modifies the acceleration of the scale factor. The imaginary nature of this term suggests that it may not significantly affect the expansion of the universe in the large wavelength regime.
These acceleration equations provide insights into the dynamics of the scale factor in the presence of matter, radiation, and phason contributions. They demonstrate how the quantum mechanical description of particles in a box with moving walls can give rise to cosmological equations that resemble the Friedmann equations of general relativity. The inclusion of the phason term in Equation (13) highlights the potential role of quasicrystalline structures and their associated quasiparticles in the evolution of the universe.
Our analysis commenced with the time-independent Schrödinger equation for a particle confined within a potential well. We then introduced several crucial modifications to the system: a general energy-balancing constraint (Equation 9), a time-dependent boundary condition representing moving walls, and a dynamically evolving volume. These additions were aimed at establishing a connection between the quantum and classical regimes across different scales. By applying the Hubble law to the wavelength of the particle and relating it to the cosmic scale factor, we derived a Friedmann-like equation (Equation 5) from the quantum mechanical description of the system.
In the relativistic extension of our analysis, we employed the Klein-Gordon equation and the relativistic energy-momentum relation (Equation 8). This led to the emergence of a new term in the resulting Friedmann-like equation (Equation 11), which bears a resemblance to the cosmological constant term in the standard Friedmann equations of general relativity.
Furthermore, we explored the implications of different dispersion relations, particularly those associated with phonons and phasons in quasicrystalline structures. The presence of an imaginary term in the phason dispersion relation (Equation 12) led to a modified Friedmann-like equation (Equation 13) that includes a dissipative contribution arising from the phason friction coefficient. This can be interpreted as having different regimes contributions.
Bridging quantum mechanics and cosmology is complex and challenging. Our simplified model offers insights but may not capture all intricacies of this transition. A comprehensive framework incorporating quantum gravity, quantum field theory, and cosmological perturbation theory might be needed to fully understand the interplay between quantum phenomena and the universe’s large-scale structure.
Nevertheless, our analysis demonstrates the potential for deriving cosmological equations from the quantum mechanical description of a particle in a box with time-dependent boundary conditions at intermediate scales. The derivation of Friedmann-like equations indicates a profound connection between the microscopic quantum world and the macroscopic evolution of the universe. This suggests that the local expansion, influenced by matter distributions, is analogous to the wave-like properties of quantum particles. The inclusion of phonon and phason contributions from quasicrystalline structures further emphasizes the potential role of condensed matter physics in understanding the nature of dark energy, dark matter, and the expansion of the universe.

### v2

@@ Abstract
We propose a heuristic model of the universe as a growing quasicrystal projected from a higher-dimensional lattice. This quasicrystalline framework offers a novel perspective on cosmic expansion, where the intrinsic growth dynamics naturally give rise to the observed large-scale expansion of the universe. Motivated by this model, we explore the Schrödinger equation for a particle in a box with time-dependent boundaries, representing the expanding underlying space. By introducing a constraint that links microscale quantum phenomena with macroscale cosmological quantities, we derive an equation resembling the Friedmann equation, providing potential insights into the Hubble tension. Our model incorporates phonons and phasons—quasiparticles inherent in quasicrystalline structures—that play critical roles in cosmic-scale dynamics and the universe’s expansion. This framework suggests that the necessity for an inflationary period may be obviated. Furthermore, phonons arising from the quasicrystalline structure may serve as dark matter candidates, influencing the dynamics of ordinary matter while remaining largely undetectable through electromagnetic interactions. Drawing parallels with crystalline matter at atomic scales, which is fundamentally quantum in nature, we explore how the notion of tiling space can support continuous symmetry atop a discrete structure. This provides a novel framework for understanding the universe’s expansion and underlying structure. Consequently, our approach suggests that further development could enhance our understanding of cosmic expansion and the universe’s structure, bridging concepts from quantum mechanics, condensed matter physics, and cosmology.
Keywords Quantum Gravity, Cosmology, Hubble Tension, Quasicrystals
@@ 1 Introduction
Recent developments in theoretical physics have propelled us toward a deeper understanding of spacetime and gravity. The field of emergent gravity [1] suggests that spacetime and gravity may be byproducts of fundamental quantum mechanical processes. This paradigm shift is supported by theories such as loop quantum gravity and spin foam [2], as well as superstring theory and M theory [3], all of which propose a quantum-level discreteness underlying gravity.
A key focus since the advent of perturbative superstring theory has been the quest to unravel quantum gravity [4]. This pursuit requires a deep understanding of the theory’s fundamental principles and extracting insights about Planck-scale physics [5]. Consequently, there is a growing trend to incorporate concepts from condensed matter or solid-state physics, such as viewing spacetime as a deformable solid or crystal [6]. More recently, quasicrystals have been investigated [7, 8].
@@
Our research adopts a dual approach within this context. First, we draw inspiration from the concept of a deformable solid or crystal and quantum cosmology, which deals with discreteness, and we propose modeling spacetime as a growing quasicrystal. Characterized by long-range order without periodicity, this structure reflects the aperiodic patterns found in condensed matter physics. Our principal result suggests that the cosmological scale factor can be interpreted as an intrinsic scaling property within a quasicrystal tiling space framework. We further explore the implications of this model for resolving the Hubble tension, offering a novel perspective on the universe’s evolution and providing insights into other unresolved issues in cosmological models, including dark matter and dark energy. A critical aspect of quasicrystal growth is that external expansion may be accompanied by internal subdivision and rescaling, which parallels the observed expansion of the universe, where space appears to be generated between galaxies.
Second, we build upon conventional and relativistic quantum mechanics. While the universe’s expansion is traditionally addressed from a classical perspective on large scales, this expansion must also extend down to microscopic scales, as suggested by various approaches in quantum cosmology [12, 13] and by our quasicrystal growth model. Although Hubble-Lemaître’s law is empirically validated at cosmological distances, we extend its conceptual framework to ‘intermediate’ and even microscopic scales as a theoretical proposition. While these scales are not typically associated with cosmic expansion, our quasicrystal model of spacetime provides a basis for such an extension. This hypothesis enables us to explore the implications of spacetime expansion on quantum states confined within matter-dominated regions.
We analyze the wave function of a particle with mass $m$ within a specific spatial domain using both the Schrödinger equation and the Klein-Gordon equation. Crucially, we posit that the wavenumber, wavelength, and quantum amplitudes become functions of the dynamic size of the spatial boundary and are explicitly subject to the Hubble-Lemaître’s law. Considerations of the evolving boundary of this spatial domain lead us to a modified version of the Friedmann equation, a key component in cosmological models, thereby highlighting the emergence of classical spacetime from quantum mechanical principles. We apply this framework to the later stages of the universe’s evolution under conventional quantum mechanics to address the global growth and rescaling of the quasicrystal structure. The early universe, where curved spacetime geometry plays a significant role, is left as a subject for future work.
@@
Our research converges on the concept of rescaling the underlying geometry along with emerging constraints relevant to physics at different scales. This approach aligns with our framework of conceptualizing spacetime as a quasicrystalline structure. By incorporating scale invariance, we aim to establish a connection between the microscale domains of quantum mechanics and the macroscale phenomena observed in cosmology. This synthesis provides a unified perspective that may offer deeper insights into the fundamental nature of the universe.
@@
In the subsequent sections, we will explore the notion of spacetime as a tiling space, drawing parallels with quasiperiodic structures in condensed matter physics to strengthen our hypothesis regarding emergent spacetime and gravity. Additionally, we will expand on the mathematical foundations of our model, examining how energy considerations translate into cosmological parameters.